OpenStudy (anonymous):

What is the least common denominator of 28 and 32?

OpenStudy (anonymous):

28=2*2*7 32=2*2*2*2*2 Least Common Denominator=2*2* 7* 2*2*2=224

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

You'r welcome